OpenStudy (anonymous):

Use the arc length formula and the given information to find r. Show your work for full credit. s = 8.1 ft θ = pi/3 rad r = ?

OpenStudy (anonymous):

@amistre64 @Michele_Laino

563blackghost (563blackghost):

We will use the formula...\[s=r \theta \] We first input what we know...

OpenStudy (anonymous):

So we have 8.1=r(pi/3)

OpenStudy (anonymous):

I am just not sure where to go from here

563blackghost (563blackghost):

Correct! So we need to now simplify se we are left with r.... so we will do this\[\frac{ 8.1 }{ \pi /3 }\]

OpenStudy (anonymous):

Okay, do we simplify the bottom so that we have a number to divide by?

OpenStudy (anonymous):

pi/3=1.04?

563blackghost (563blackghost):

Yup! And then what you get you will divide by 8.1 to get r.

OpenStudy (anonymous):

Okay! So I got r=7.78

563blackghost (563blackghost):

Yup! I would use a calculator to be able to find the exact simplification of pi divided by 3 because if you round you will get 7.78 which close but the answer is 7.73493...

OpenStudy (anonymous):

Okay thank you so much!

563blackghost (563blackghost):

Np xD Glad to help! "Thank You" for asking the question! xD